# Characterizing Pattern Matching and Its Limits on Compositional Task Structures

### v2

@@ Appendix C Implementation Details for the Coverage Determination Algorithm
Input: Training examples $D=\{(x_{i},f(x_{i}))\}_{i=1}^{N}$, where $x_{i}\in{\mathcal{X}}^{n}$ and $f(x_{i})\in{\mathcal{X}}$
Minimum evidence threshold $k\geq 1$
Output: Coverage set $\text{Cover}(D)$
@@
foreach nonempty subset $I\subsetneq[n]$ do
 $\text{Behavior}_{I}\leftarrow$ map from subsequence $x_{I}$ to the mapping $\{x_{[n]\setminus I}\mapsto f(x)\mid x\in D\}$
end foreach
/* STEP 2: Identify functionally equivalent subsequences */
foreach nonempty subset $I\subsetneq[n]$ do
 $UF_{I}\leftarrow\text{new UnionFind}()$
 foreach pair of subsequences $(\alpha,\beta)$ in $\text{Behavior}_{I}$ do
    $\text{SharedComplements}\leftarrow$ complements observed with both $\alpha$ and $\beta$
    if No contradictions in SharedComplements and matching evidence $\geq k$ then
       $UF_{I}.\text{Union}(\alpha,\beta)$ ;
       // Mark as functionally equivalent
      end if
   end foreach
 $\text{EquivClasses}_{I}\leftarrow UF_{I}$
end foreach
/* STEP 3: Build substitution graph */
$G\leftarrow\text{empty graph}$ with nodes for all $x\in{\mathcal{X}}^{n}$
foreach pair of inputs $(x,y)$ with $f(x)=f(y)$ do
 foreach subset II where $x$ and $y$ differ only on indices in II do
    if $\text{EquivClasses}_{I}.\text{Find}(x_{I})=\text{EquivClasses}_{I}.\text{Find}(y_{I})$ then
         Add edge $(x,y)$ to $G$
       break
      end if
   end foreach
end foreach
/* STEP 4: Determine coverage */
$\text{Cover}(D)\leftarrow\bigcup_{x\in D}$ ConnectedComponent(G,x)(G,x)
return $\text{Cover}(D)$
@@
Alg.˜1 presents our approach to computing the coverage set with a minimum evidence threshold $k$. The algorithm works in four main stages:
Stage 1: Behavior mapping
We first analyze the training data to create a mapping of behaviors for each possible subsequence of the input. For each subset of indices $I$, we record how different subsequences $x_{I}$ behave when paired with their complements $x_{[n]\setminus I}$, essentially mapping each subsequence to a function from complements to outputs.
Stage 2: Equivalence class construction
For each subset of indices $I$, we build equivalence classes of subsequences that exhibit functionally identical behavior. Two subsequences are considered equivalent only if: (1) they share at least $k$ distinct complements where they produce the same output, and (2) they never produce different outputs when given the same complement (no contradictions). We use a Union-Find data structure to efficiently track and merge these equivalence classes. The Union-Find (or Disjoint-Set) data structure efficiently maintains a collection of disjoint sets, supporting two key operations: (1) Find - determine which set an element belongs to, and (2) Union - merge two sets.
Stage 3: Substitution Graph Construction
We construct a graph where nodes represent input sequences from our training and test sets, rather than the entire domain space (which would be computationally prohibitive for large token sets). We add an edge between two inputs $x$ and $y$ if and only if: (1) they produce the same output, (2) they differ only in one subsequence position set $I$, and (3) their differing subsequences belong to the same equivalence class. This graph represents the space of safe substitutions where one can replace a subsequence with a functionally equivalent alternative without changing the expected output. Our implementation uses parallel processing to efficiently construct this graph, even for large datasets.
Stage 4: Coverage computation
Finally, we compute the coverage set by taking the union of all connected components in the substitution graph that contain at least one training example. This set comprises all inputs that are reachable from the training data through chains of equivalent subsequence substitutions.
@@ D.1 Causal Tracing Methodology
To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference (goldowsky2023localizing; hanna2023does). Specifically, we measure the causal effect using the indirect effect metric defined in (sharma2024locating). This methodology allows us to identify which components and positions in the model most strongly contribute to compositional generalization. We illustrate the measurement with 2-Hop task.
@@
$\text{IE}_{h_{i}^{(\ell)}}=\frac{p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})-p(t^{*})}{p^{*}(t^{*})-p(t^{*})}$ (4)
@@
  • $p(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the clean run $G$
  • $p^{*}(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the corrupted run $G^{*}$
  • $p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the patched run $G[\leftarrow{h^{*}}_{i}^{(\ell)}]$
@@ D.2 Causal Tracing Results for Each $k$-Cutoff Value in 2-Hop Task
Figure 8 displays the causal tracing results for the 2-Hop task, broken down by different $k$-cutoff values. We observe that the causal patterns are similar across different $k$-cutoff values, with slight differences in where and how strongly the causal effects manifest in the model. This suggests that once an example falls within coverage (even with minimal evidence, $k=1$), the model forms internal representations that play similar causal roles in prediction.
@@ D.3 Token Set Size Ablation
We show that the observed patterns of cosine similarity analysis and causal tracing in the 2-Hop task are consistent across different token set sizes $|{\mathcal{X}}|$. For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99. Figure Fig.˜9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases. The causal tracing results in Fig.˜10 show that the clustered functional equivalence representations at the lower layers of position $x_{2}$ play a causal role in determining the model’s final prediction.
@@ D.4 Task Ablation
We show that GPT-2 models trained on Parallel-2-Hop and 3-Hop tasks exhibit the same patterns: clustered functional equivalence representations of intermediate states at specific layers and positions, confirmed through cosine similarity analysis, with causal tracing analysis verifying their role in model predictions. For both tasks, we analyze with $|{\mathcal{X}}|=50$ and examine model checkpoints with training dataset size $N=\hat{N}_{\text{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
Figures 11 and 12 show the results for the Parallel-2-Hop task. The IICG patterns reveal strong representation clustering at mid-layers: at positions $x_{2}$ and $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{4}$ when grouped by $b_{2}=f_{2}(x_{3},x_{4})$. Causal tracing confirms the causal role of these clustered representations in the model’s predictions.
@@
Similarly, Figures 13 and 14 show results for the 3-Hop task. The IICG patterns exhibit strong representation clustering at mid-layers: at position $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{3}$ when grouped by $b_{2}=f_{2}(b_{1},x_{3})$. Causal tracing again confirms the causal importance of these representations.
@@ Appendix F Proofs: Asymptotically Perfect Coverage in 2-Hop Task
  • In Sec.˜F.3, we provide (almost) all postponed proofs of lemmas used in Sec.˜F.2. We first prove that the connectivity of all $b$-evidence graphs (Def.˜F.5), which are the intermediate-state-specific induced subgraphs of the substitution graph, is a sufficient condition for the perfect in-domain generalization we want to show: see Lemma˜F.10.
@@ F.1 Preliminaries: Definitions, Assumptions, and Facts
Given a train dataset $D$ and for each ${\bm{x}}\in{\mathbb{X}}$, define $Z_{\bm{x}}$ as the sampled count of $\bm{x}$} (i.e., the number of identical duplicates equal to $\bm{x}$}) in $D$. Then, if $D$ is generated by with-replacement sampling $N$ times from $\mathbb{X}$}, the random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ follows a multinomial distribution with parameters $n=N$ and $p_{\bm{x}}=\frac{1}{\left|{\mathbb{X}}\right|}$ ($\forall{\bm{x}}\in{\mathbb{X}}$). That is, for all collections of integers $z_{\bm{x}}\geq 0$ ($\forall{\bm{x}}\in{\mathbb{X}}$) such that $\sum_{{\bm{x}}\in{\mathbb{X}}}z_{\bm{x}}=N$,
@@ F.2 Main Theorem: Sample Complexity Bound for Perfect Coverage With High Probability
Lemma F.12 (De-Poissonizaiton of Monotone Multinomial Events, $\downarrow$).
@@
Lemma F.14 (zhao2014k, Theorem 2; zhao2017connectivity, Theorem 1 & Remark 1).
@@
Then, choosing all $n_{b}=\left|V_{b}\right|$ to be greater than $n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right)$, we yield the bound (9). Note that a choice of $p$ larger than the threshold in Eq.˜8 will never change that ${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is connected})\to 1$ as $n\to\infty$. Thus, since we choose $\alpha_{n}=\ln n$, it suffices to have
@@
Even for $k=1$, we follow almost the same proof as that of Theorem˜F.7, except for the last few steps. Namely, we again observe that every ${\mathcal{G}}_{b}^{(D,1)}$ is a binomial random $1$-intersection graph ${\mathcal{G}}^{(1)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-(N-{\varepsilon})/\left|{\mathbb{X}}\right|)$ (see the definition of binomial random intersection graph in Sec.˜F.3.2). To guarantee the union bound in Eq.˜9, however, we use a different lemma: Lemma˜F.15 (instead of Lemma˜F.14). In particular, we will use only the $\rho\leq 1$” part (which tightly matches our assumption $m=O(n_{b})$ ($b\in B_{D}$)) of the lemma below:
Lemma F.15 (singer1995random, Propositions 3.1–2, Theorem 3.3).
@@
Hence, to show Eq.˜10, it suffices to prove the following three inequalities:
@@
A size $N$ of training dataset $D$ that is sufficient to ensure Eqs.˜13 and 14 can be characterized with the following lemma.
@@
is enough to ensure both Eqs.˜13 and 14. The remaining task now is to determine a condition for $N$ that guarantees Eq.˜12.
Observe that ${\mathcal{M}}_{1}$ (defined as Eq.˜11) is a non-decreasing property (Def.˜F.11) because of a similar reason for the monotonicity of the property $\mathcal{C}$} (defined as Eq.˜7; see the proof of Theorem˜F.7). Thanks to the monotonicity of ${\mathcal{M}}_{1}$, we can apply the Poissonization technique (Lemma˜F.12) once again. That is, taking any fixed $c\in(0,2)$, we have
@@
for any ${\varepsilon}\in(0,cN)$, where ${\mathbb{P}}_{\mathrm{Po}(N+{\varepsilon})}$ is the probability measure under i.i.d. Poisson random variables $Z_{\bm{x}}\sim\mathrm{Poisson}\left(\frac{N+{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)$ ($\forall{\bm{x}}\in{\mathbb{X}}$). Let us take ${\varepsilon}=\sqrt{\frac{6N}{3-c}\ln\left(\frac{4}{\delta}\right)}$, which is smaller than $N$ by Eq.˜15 (since $\left|{\mathcal{X}}_{12}\right|\geq 6>\frac{6}{3-c}$) and ensures that
@@
Then, choosing $\hat{n}$ greater than $n_{0}\left(\frac{\delta}{4}\right)$, we yield the bound (16). Observe that a choice of $p$ smaller than the threshold in Eq.˜8 will never change the limit ${\mathbb{P}}(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1)\to 0$. Thus, by our choice $\alpha_{n}=-(1-\frac{1}{k})\ln n$, it is enough to have
@@
Before moving on to the postponed proofs of lemmas, we lastly remark that the same proof (of Theorem˜F.16) can hardly apply to the case of $k=1$ in general. This is because, for the sake of simplicity in applying the necessary condition for a perfect coverage (Lemma˜F.18), we first characterize a minimal data size to guarantee $\overline{D}={\mathbb{X}}$ with high probability as Eq.˜15, using the tail bound of the coupon collector’s problem (Lemma˜F.19). Unfortunately, it already exceeds the sample complexity upper bound to ensure $\overline{D}\subset\mathrm{Cover}_{1}(D)$ (obtained as Theorem˜F.8) for $k=1$ with high probability, especially in the regime of balanced cardinalities assumed in Corollary˜F.17.

### v3

@@ Appendix C Implementation Details for the Coverage Determination Algorithm
In this section, we explain our algorithm of computing the $k$-coverage $\mathrm{Cover}_{k}(D)$ (see Def.˜3.2), given a train dataset $D\subseteq{\mathcal{X}}^{\ell}$ and a threshold $k$ to check functional equivalences (see Def.˜3.1). The algorithm works in three main stages, as described with a pseudo-code in Alg.˜1.
[…]
Output: Coverage set: $\mathrm{Cover}_{k}(D)$
1
@@
2 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav$$\leftarrow$ an empty hash table
3foreach $(I,I^{\mathsf{c}}):\varnothing\neq I\subsetneq[\ell],~I^{\mathsf{c}}=[\ell]\setminus I$ do
4
5 ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I]\leftarrow$ an empty hash table
6 foreach ${\bm{x}}\in D$ do
7    if ${\bm{x}}_{I}\notin{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I].keys$ then
8       ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{x}}_{I}]\leftarrow$ an empty hash table
9      end if
10    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{x}}_{I}][{\bm{x}}_{I^{\mathsf{c}}}]\leftarrow f({\bm{x}})$
11   end foreach
12
13 end foreach
14
/* STEP 2: Build substitution graph using behavior maps */
15 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}$$\leftarrow$ a graph (vertex set $\mathcal{X}$$^{\ell}$; edge set $\varnothing$)
16foreach $(I,I^{\mathsf{c}}):\varnothing\neq I\subsetneq[\ell],~I^{\mathsf{c}}=[\ell]\setminus I$ do
17 foreach $({\bm{a}},{\bm{a}}^{\prime}):{\bm{a}},{\bm{a}}^{\prime}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I].keys$  and  ${\bm{a}}\neq{\bm{a}}^{\prime}$ do
    ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc}\leftarrow{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{a}}].keys~\cap~{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{a}}^{\prime}].keys$ // II-co-occurrences
18
19    if $|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc}|\geq k$  and  $\forall{\bm{b}}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc}:{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{a}}][{\bm{b}}]={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav}[I][{\bm{a}}^{\prime}][{\bm{b}}]$   then
       /* Statisfies the functional $k$-equivalence (Def.˜3.1) */
20       foreach ${\bm{b}}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc}$ do
21            Let ${\bm{x}},{\bm{x}}^{\prime}\in{\mathcal{X}}^{\ell}$ such that ${\bm{x}}_{I}={\bm{a}},~{\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime},$ and ${\bm{x}}_{I^{\mathsf{c}}}={\bm{x}}^{\prime}_{I^{\mathsf{c}}}={\bm{b}}$
22           Add an edge $\{{\bm{x}},{\bm{x}}^{\prime}\}$ to $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}$}
23         end foreach
24
25      end if
26
27   end foreach
28
29 end foreach
30
/* STEP 3: Build $k$-coverage */
31 $\mathrm{Cover}_{k}(D)\leftarrow\bigcup_{{\bm{x}}\in D}\mathrm{ConnectedComponent}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}};{\bm{x}})$
return $\mathrm{Cover}_{k}(D)$
@@
In the first stage (Alg.˜1, Lines 11), we first create a mapping of behaviors for all subsequences of every training data. To elaborate on the behaviors, for each subset of indices $I$, we record how each subsequence ${\bm{x}}_{I}:=(x_{i})_{i\in I}$ at $I$ of ${\bm{x}}\in D$ behave when being paired with another subsequence ${\bm{x}}_{I^{\mathsf{c}}}$ at the complement of indices $I^{\mathsf{c}}:=[\ell]\setminus I$. This ends up creating a thrice-nested mapping $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav$}, as showcased in Line 1. It is particularly useful in the next stage, where we should collect the co-occurrences of each pair of input sequences and verify their consistencies.
The second stage (Alg.˜1, Lines 11) is for establishing the substitution graph ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}}={\mathcal{G}}^{(D,k)}$ which is essential in the definition of $k$-coverage. To recall, ${\mathcal{G}}^{(D,k)}$ is an undirected graph with vertex set $\mathcal{X}$$^{\ell}$, having edges $\{{\bm{x}},{\bm{x}}^{\prime}\}$ when and only when there exists an index set $I$ such that ${\bm{x}}_{I}\equiv^{(D,k)}_{I}{\bm{x}}^{\prime}_{I}$. To verify the presence of an edge via functional $k$-equivalence, we need to figure out at least $k$ distinct $I$-co-occurrences ($\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt CoOcc$} in the pseudocode) for some index set $I$ (Def.˜3.1(1)) and check whether all $I$-co-occurrences satisfy the consistency (Def.˜3.1(2)). It is done for every nonempty index set $I\subsetneq[\ell]$ and every distinct pair of subsequences $({\bm{a}},{\bm{a}}^{\prime})\in{\mathcal{X}}^{|I|}\times{\mathcal{X}}^{|I|}$, where we take advantage of the mapping $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\tt Behav$} defined in stage 1; see Line 1 in particular.
In the last stage, we generate the $k$-coverage (Alg.˜1, Line 1). Here we utilize a subroutine ${\rm ConnectedComponent}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}};{\bm{x}})$ which returns the set of all vertices of $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}$} sitting on the same connected component as the given vertex $\bm{x}$}. Hence, Line 1 generates the set of all input sequences that are connected with a train datapoint by a path in the substitution graph $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}{\mathcal{G}}$}, which is precisely the definition of $k$-coverage. In other words, this set comprises all inputs that are reachable from the training data through a chain of functionally equivalent subsequence substitutions.
@@ D.1 Causal Tracing Methodology
To analyze the causal role of specific hidden representations in our Transformer model, we employ causal tracing, a technique that measures the effect of intervening on intermediate activations during inference (Goldowsky-Dill et al., 2023; Hanna et al., 2023). Specifically, we measure the causal effect using the indirect effect metric defined in (Sharma et al., 2024). This methodology enables us to identify which components and positions in the model most strongly contribute to compositional generalization. We illustrate the measurement with the 2-Hop task.
@@
$\mathrm{IE}_{h_{i}^{(\ell)}}=\frac{p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})-p(t^{*})}{p^{*}(t^{*})-p(t^{*})}$
@@
  • $p(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the clean run $G$;
  • $p^{*}(t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the corrupted run $G^{*}$;
  • $p[\leftarrow{h^{*}}_{i}^{(\ell)}](t^{*})$ is the probability assigned to the corrupted output $t^{*}$ in the patched run $G[\leftarrow{h^{*}}_{i}^{(\ell)}]$.
@@ D.2 Causal Tracing Results for Each $k$-Cutoff Value in 2-Hop Task
Fig.˜8 displays the causal tracing results for the 2-Hop task, broken down by different $k$-cutoff values. We observe that the causal patterns are similar across different $k$-cutoff values, with slight differences in where and how strongly the causal effects manifest in the model. This suggests that once an example falls within coverage (even with minimal evidence, $k=1$), the model forms internal representations that play similar causal roles in prediction.
@@ D.3 Token Set Size Ablation
We show that the observed patterns of cosine similarity analysis and causal tracing in the 2-Hop task are consistent across different token set sizes $|{\mathcal{X}}|$. For $|{\mathcal{X}}|=70,100,150,200$, we analyze model checkpoints with training dataset size $N=\hat{N}_{\mathrm{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99. Figure Fig.˜9 shows the results, indicating strong representation clustering at the lower layers of position $x_{2}$ for all cases. The causal tracing results in Fig.˜10 show that the clustered functional equivalence representations at the lower layers of position $x_{2}$ play a causal role in determining the model’s final prediction.
@@ D.4 Task Ablation
We show that GPT-2 models trained on Parallel-2-Hop and 3-Hop tasks exhibit the same patterns: clustered functional equivalence representations of intermediate states at specific layers and positions, confirmed through cosine similarity analysis, with causal tracing analysis verifying their role in model predictions. For both tasks, we analyze with $|{\mathcal{X}}|=50$ and examine model checkpoints with training dataset size $N=\hat{N}_{\mathrm{req}}(|{\mathcal{X}}|)$ that achieve training accuracy > 0.99.
Figs.˜11 and 12 show the results for the Parallel-2-Hop task. The IICG patterns reveal strong representation clustering at mid-layers: at positions $x_{2}$ and $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{4}$ when grouped by $b_{2}=f_{2}(x_{3},x_{4})$. Causal tracing confirms the causal role of these clustered representations in the model’s predictions.
@@
Similarly, Figs.˜13 and 14 show results for the 3-Hop task. The IICG patterns exhibit strong representation clustering at mid-layers: at position $x_{3}$ when grouped by $b_{1}=f_{1}(x_{1},x_{2})$, and at position $x_{3}$ when grouped by $b_{2}=f_{2}(b_{1},x_{3})$. Causal tracing again confirms the causal importance of these representations.
@@ Appendix F Proofs: Asymptotically Perfect Coverage in 2-Hop Task
  • In Sec.˜F.3, we provide (almost) all postponed proofs of lemmas used in Sec.˜F.2. We first prove that the connectivity of all $b$-evidence graphs (Def.˜F.5), the intermediate-state-specific induced subgraphs of the substitution graph, is a sufficient condition for the perfect in-domain generalization we want to show: see Lemma˜F.10.
@@ F.1 Preliminaries: Definitions, Assumptions, and Facts
Given a train dataset $D$ and for each ${\bm{x}}\in{\mathbb{X}}$, define $Z_{\bm{x}}$ as the sampled count of $\bm{x}$} (i.e., the number of identical duplicates equal to $\bm{x}$}) in $D$. In particular, if $D$ is uniformly randomly generated by with-replacement sampling $N$ times from $\mathbb{X}$}, the random vector $(Z_{\bm{x}})_{{\bm{x}}\in{\mathbb{X}}}$ follows a multinomial distribution with parameters $n=N$ and $p_{\bm{x}}=\frac{1}{\left|{\mathbb{X}}\right|}$ ($\forall{\bm{x}}\in{\mathbb{X}}$). That is, for all collections of integers $z_{\bm{x}}\geq 0$ ($\forall{\bm{x}}\in{\mathbb{X}}$) such that $\sum_{{\bm{x}}\in{\mathbb{X}}}z_{\bm{x}}=N$,
@@ F.2 Main Theorem: Sample Complexity Bound for Perfect Coverage With High Probability
Lemma F.12 (De-Poissonization of Monotone Multinomial Events, $\downarrow$).
@@
Lemma F.14 (Zhao et al., 2014, Theorem 2; Zhao et al., 2017, Theorem 1 & Remark 1).
@@
Then, choosing all $n_{b}=\left|V_{b}\right|$ to be greater than $n_{0}\!\left(\frac{\delta}{2\left|B_{D}\right|}\right)$, we yield the bound (8). Note that a choice of $p$ larger than the threshold in Eq.˜7 will never change that ${\mathbb{P}}({\mathcal{G}}^{(k)}(n,m,p)\text{ is connected})\to 1$ as $n\to\infty$. Thus, since we choose $\alpha_{n}=\ln n$, it suffices to have
@@
Even for $k=1$, we follow almost the same proof as that of Theorem˜F.7, except for the last few steps. Namely, we again observe that every ${\mathcal{G}}_{b}^{(D,1)}$ is a binomial random $1$-intersection graph ${\mathcal{G}}^{(1)}(n_{b},m,p)$ with parameters $n_{b}=\left|V_{b}\right|$, $m=\left|{\mathcal{X}}_{3}\right|$, and $p=1-\exp(-(N-{\varepsilon})/\left|{\mathbb{X}}\right|)$ (see the definition of binomial random intersection graph in Sec.˜F.3.2). To guarantee the union bound in Eq.˜8, however, we use a different lemma: Lemma˜F.15 (instead of Lemma˜F.14). In particular, we will use only the $\rho\leq 1$” part (which tightly matches our assumption $m=O(n_{b})$ ($b\in B_{D}$)) of the lemma below:
Lemma F.15 (B., 1995, Propositions 3.1–2, Theorem 3.3).
@@
Hence, to show Eq.˜9, it suffices to prove the following three inequalities:
@@
A size $N$ of training dataset $D$ that is sufficient to ensure Eqs.˜12 and 13 can be characterized with the following lemma.
@@
is enough to ensure both Eqs.˜12 and 13. The remaining task now is to determine a condition for $N$ that guarantees Eq.˜11.
Observe that ${\mathcal{M}}_{1}$ (defined as Eq.˜10) is a non-decreasing property (Def.˜F.11) because of a similar reason for the monotonicity of the property $\mathcal{C}$} (defined as Eq.˜6; see the proof of Theorem˜F.7). Thanks to the monotonicity of ${\mathcal{M}}_{1}$, we can apply the Poissonization technique (Lemma˜F.12) once again. That is, taking any fixed $c\in(0,2)$, we have
@@
for any ${\varepsilon}\in(0,cN)$, where ${\mathbb{P}}_{\mathrm{Po}(N+{\varepsilon})}$ is the probability measure under i.i.d. Poisson random variables $Z_{\bm{x}}\sim\mathrm{Poisson}\left(\frac{N+{\varepsilon}}{\left|{\mathbb{X}}\right|}\right)$ ($\forall{\bm{x}}\in{\mathbb{X}}$). Let us take ${\varepsilon}=\sqrt{\frac{6N}{3-c}\ln\left(\frac{4}{\delta}\right)}$, which is smaller than $N$ by Eq.˜14 (since $\left|{\mathcal{X}}_{12}\right|\geq 6>\frac{6}{3-c}$) and ensures that
@@
Then, choosing $\hat{n}$ greater than $n_{0}\left(\frac{\delta}{4}\right)$, we yield the bound (15). Observe that a choice of $p$ smaller than the threshold in Eq.˜7 will never change the limit ${\mathbb{P}}(\min\deg{\mathcal{G}}^{(k)}(n,m,p)\geq 1)\to 0$. Thus, by our choice $\alpha_{n}=-(1-\frac{1}{k})\ln n$, it is enough to have
@@
Before moving on to the postponed proofs of lemmas, we lastly remark that the same proof (of Theorem˜F.16) can hardly apply to the case of $k=1$ in general. This is because, for the sake of simplicity in applying the necessary condition for a perfect coverage (Lemma˜F.18), we first characterize a minimal data size to guarantee $\overline{D}={\mathbb{X}}$ with high probability as Eq.˜14, using the tail bound of the coupon collector’s problem (Lemma˜F.19). Unfortunately, it already exceeds the sample complexity upper bound to ensure $\overline{D}\subset\mathrm{Cover}_{1}(D)$ (obtained as Theorem˜F.8) for $k=1$ with high probability, especially in the regime of balanced cardinalities assumed in Corollary˜F.17.